# Peierls substitution and Hall motion in exotic Carroll dynamics

### v1

@@ IV Exotic Carroll particle
$B=0\,,$ (IV.13)
then, assuming that $\,\theta\,\kappa_{mag}\neq 1,$ the position satisfies
$\dot{x}_{i}=-\frac{e\theta}{1-\theta\kappa_{mag}}\epsilon_{ij}E_{j}\,.$ (IV.14)
Thus when $\theta\neq 0$ Hall motions persist even for $\kappa_{mag}=0$. (IV.14) is reminiscent of “the” anomalous Hall effect observed in ferromagnetic materials in the absence of an external magnetic field Kar+ ; Ezawa ; Stone ; AnomHall . Comparison of (IV.14) with (IV.7) underlines that $\kappa_{mag}$ behaves indeed as a proper magnetic field carried by the particle, and could be viewed as the planar analog of a magnetic monopole. It is also reminiscent of the rather mysterious magnetic field of relativistic anyons FeherAny .
We underline that the system is regular in both of these particular cases (IV.11) and (IV.13).
IV.1 The singular Carroll case
In the singular case (IV.9) both equations in (IV.5) are contradictory unless $e{\mathbf{E}}=0$, when they are undetermined. This is unlike as in exotic Galilean mechanics, eqn. (III.2), where the $p$-terms on the rhs can compensate the exotic one, allowing, for vanishing effective mass, motions which follow the Hall law, (III.11) DHPeierls ; DHJPA ; NCLandau ; ExoRev ; ZHChiral .
We resort to Hamiltonian reduction FaTMF ; FaJa ; DHJPA ; NCLandau along the lines as for exotic Galilei DHJPA ; NCLandau . We start with the regular case $m^{*}\neq 0$ and introduce new, canonical coordinates DHJPA ,
$Q_{i}$ $=x_{i}+\frac{1}{B^{*}}\left[1-\sqrt{\frac{m^{*}}{m}}\right]\,\epsilon_{ij}p_{j}\,,$ (IV.15a)
$P_{i}$ $=\sqrt{\frac{m^{*}}{m}}\,p_{i}-\hbox{$\textstyle\frac{1}{2}$}B^{*}\epsilon_{ij}Q_{j}\,,$ (IV.15b)
in terms of which the symplectic form and Hamiltonian are,
$\Omega_{exo}$ $=dP_{i}\wedge dQ_{i}\,,$ (IV.16a)
${\mathscr{H}}_{exoC}$ $=-\frac{1}{2}\left(\sqrt{\frac{m}{m^{*}}}+1\right)eE_{i}dQ_{i}+\frac{1}{B^{*}}\left(\sqrt{\frac{m}{m^{*}}}-1\right)\epsilon_{ij}eE_{i}dP_{j}\;.$ (IV.16b)
The corresponding Poisson brackets and equations of motion are
$\left\{Q_{i},P_{j}\right\}=\delta_{ij},\text{ \ }\left\{Q_{i},Q_{j}\right\}=0,\text{ \ }\left\{P_{i},P_{j}\right\}=0$ (IV.17)
and
$\dot{P}_{i}=\frac{1}{2}\left(\sqrt{\frac{m}{m^{*}}}+1\right)eE_{i}{\quad\text{and}\quad}\dot{Q}_{i}=-\frac{1}{B^{*}}\left(\sqrt{\frac{m}{m^{*}}}-1\right)\epsilon_{ij}\,eE_{j}\,.$ (IV.18)
respectively. In the critical case (IV.9) the inverse transformation,
$x_{i}=\frac{1}{2}\left(\sqrt{\frac{m}{m^{*}}}+1\right)Q_{i}-\frac{1}{B^{*}}\left(\sqrt{\frac{m}{m^{*}}}-1\right)\epsilon_{ij}P_{j}\,,$ (IV.19)
becomes singular (IV.9), the velocities diverge, however the coordinates (IV.15) do have a well-defined limit. Letting $m^{*}\to 0$, (IV.15b) implies that ${\rm{\bf P}}$ is rotated rigidly by $\pi/2$ w.r.t. ${\bf Q}$,
$P_{i}=-\hbox{$\textstyle\frac{1}{2}$}B_{crit}^{*}\epsilon_{ij}Q_{j}\,,$ (IV.20)
leaving us with the Carroll guiding centre,
$Q_{i}=x_{i}+\frac{1}{B^{*}}\,\epsilon_{ij}p_{j}\,,$ (IV.21)
as only dynamical degree of freedom. The reduced symplectic form, Poisson brackets and reduced Hamiltonian are,
$\Omega=\frac{1}{2}B^{*}_{crit}\epsilon_{ij}dQ_{i}\wedge dQ_{j}\,,\qquad\big{\{}Q_{1},Q_{2}\big{\}}=-\frac{1}{B^{*}_{crit}}\,,\qquad{\mathscr{H}}_{red}=-eE_{i}Q_{i}\,$ (IV.22)
respectively, cf. (III.12). It follows that :
@@ IV.1 The singular Carroll case
In the critical case (IV.9) the guiding centre $Q_{i}$ follows the Hall law (III.16),
$\dot{Q}_{i}=\epsilon_{ij}\frac{E_{j}}{B^{*}_{crit}}\,,$ (IV.23)
whereas $P_{i}$ is determined algebraically by $Q_{i}$ as in (IV.20).
IV.2 Chiral decomposition of exotic Carroll dynamics
Further insight is provided by chiral decomposition proposed in PlyushChiral1 ; PlyushChiral2 . Starting with the exotic Souriau form $\sigma_{exoC}$ (IV.2) (or symplectic form $\Omega_{exo}$ (III.3b) and Hamiltonian, ${\mathscr{H}}_{Car}$ (IV.3)), we introduce chiral coordinates,
$p_{i}=B^{*}\epsilon_{ij}X^{-}_{j}\,,\qquad x_{i}=X^{+}_{i}+X^{-}_{i}\,.$ (IV.24)
Here $X^{+}_{i}=Q_{i}$ is in fact the guiding centre, (IV.21). The symplectic form and Hamiltonian are decomposed as,
$\Omega=\,$ $\Omega^{+}+\Omega^{-}\;\;\;=\Big{\{}\frac{B^{*}}{2}\epsilon_{ij}dX^{+}_{i}\wedge dX^{+}_{j}\Big{\}}-\Big{\{}\frac{B^{*}}{2}(1-{\theta}B^{*})\epsilon_{ij}dX^{-}_{i}\wedge dX^{-}_{j}\Big{\}},$ (IV.25a)
${\mathscr{H}}=\,$ ${\mathscr{H}}^{+}+{\mathscr{H}}^{-}=-eE_{i}X^{+}_{i}-eE_{i}X^{-}_{i}\,.$ (IV.25b)
We note the relative minus in (IV.25a) but not in (IV.25b).
$\left\{X^{+}_{i},X^{+}_{j}\right\}=-\frac{1}{B^{*}}\epsilon_{ij},\text{ \ \ }\left\{X^{-}_{i},X^{+}_{j}\right\}=0,\text{ \ \ }\left\{X^{-}_{i},X^{-}_{j}\right\}=\frac{1}{B^{*}\left(1-{\theta}B^{*}\right)}\epsilon_{ij}\,,$ (IV.26)
yield the Hamilton equations,
$\frac{dX^{+}_{i}}{ds}=\quad\;\epsilon_{ij}\frac{eE_{j}}{B^{*}}\,,\quad$ (IV.27a)
$\frac{dX^{-}_{i}}{ds}=\;-\frac{1}{\left(1-\theta{B^{*}}\right)}\epsilon_{ij}\frac{eE_{j}}{B^{*}}\,.$ (IV.27b)
which imply :
@@ IV.2 Chiral decomposition of exotic Carroll dynamics
Both chiral coordinates follow a Hall law — the guiding centre, $X^{+}_{i}$, the usual one and $X^{-}_{i}$ an anomalous one which involves also the exotic parameter $\theta$.
For $m^{*}>0$ $X^{-}_{i}$ moves in the opposite direction as $X^{+}_{i}$ does but switches direction beyond the critical value. For comparison, remember that in the Galilean case $X^{-}_{i}$ circles around the guiding centre $X^{+}_{i}$ ZHChiral . Combining $X^{+}_{i}$ and $X^{-}_{i}$ the anomalous Hall dynamics (IV.7) is recovered. Eqn. (IV.27) allows us to deduce :
Proposition IV.6.
When the non-commutative parameter is switched off then $X^{-}_{i}$ cancels the motion of the guiding centre $X^{+}_{i}$ and the particle’s position coordinate, ${\bm{x}}$, stops moving,
$\theta=0{\quad\Rightarrow\quad}\dot{{\bm{x}}}=\dot{X}^{+}+\dot{X}^{-}=0\,.$ (IV.28)
In other regular cases with $\theta\neq 0$ the motion is a combination of Hall motion of $X^{+}_{i}$ deviated by that of $X^{-}_{i}$, as illustrated in FIG.2.
The same conclusion could have been derived directly by combining ${\bm{x}}$ and ${\bm{p}}$ into
${\widetilde{Q}}_{i}=x_{i}+\theta\epsilon_{ij}p_{j}\,$ (IV.29)
which is the guiding centre ${\bf Q}$ in the critical case (IV.9), for which the equations (IV.5) imply, even for unrelated $B^{*}$ and $\theta$,
$\dot{\widetilde{Q}}_{i}=0{\quad\Rightarrow\quad}\widetilde{Q}_{i}=\mathop{\rm const.}\nolimits$ (IV.30)
Coordinate- immobility (IV.28) is recovered. when the exotic parameter is turned off, $\theta\to 0$.
We briefly recall for comparison the discussion of the exotic Galilean model in chiral terms ZHChiral . The symplectic form and thus commutation relations are still (IV.25a) and (IV.26) (up to $B^{*}\leadsto eB$), but the Hamiltonian has a kinetic term which comes from ${\bm{p}}^{2}/2m$,
${\mathscr{H}}_{exoG}=\Big{\{}-e{\mathbf{E}}\cdot{\bm{X}}^{+}\Big{\}}\;+\;\Big{\{}\dfrac{e^{2}B^{2}}{2m}({\bm{X}}^{-})^{2}-e{\mathbf{E}}\cdot{\bm{X}}^{-}\Big{\}}\,.$ (IV.31)
The system splits into two uncoupled parts, however the $X_{i}^{-}$ equation picks up a kinetic term, cf. (IV.27).
$\begin{array}[]{rlr}\displaystyle\frac{dX^{+}_{i}}{dt}&=&\;\;\epsilon_{ij}\displaystyle\frac{E_{j}}{B}\,,\\[14.0pt] \left(1-\theta{eB}\right)\,\displaystyle\frac{dX^{-}_{i}}{dt}&=&\;eB\epsilon_{ij}X^{-}_{j}-\epsilon_{ij}\displaystyle\frac{E_{j}}{B}\,.\end{array}$ (IV.32)
For $m^{*}\neq 0$ the guiding centre follows the Hall law and $X_{i}^{-}$ rotates around it. In the critical case $m^{*}=0$, $X_{i}^{+}$ moves by following the Hall law and $X_{i}^{-}$ is frozen into $X_{i}^{-}=E_{i}/eB^{2}$.
For $m^{*}=0$ both systems become singular. For Galilei it is the frequency of the rotation wich diverges, FIG.1, and for Carroll it is the velocity which becomes infinite, FIG.2.
The Souriau form of the exotic Carroll system (IV.25) splits into two parts,
$\sigma_{exoC}=\hskip 130.88268pt$ (IV.33)
[…]
cf. (III.15) in the Galilean case. The first brace rules the motion of the guiding center, and the second that of the uncoupled momentum — for any values of $B^{*}$ and $\theta$. In the singular case $m^{*}=m(1-{\theta}B^{*})=0$ (IV.33) projects, after removing the $dp_{i}\wedge ds$ term which does not contribute, to
$\sigma^{c}_{exoC}=\hbox{$\textstyle\frac{1}{2}$}B^{*}\epsilon_{ij}\,dQ_{i}\wedge dQ_{j}+eE_{i}\,dQ_{i}\wedge ds$ (IV.34)
the same as (III.18) (up to $eB\leftrightarrow B^{*}$ and $t\leftrightarrow s$), while the second braces are somewhat different. The trajectories are in general different: spiralling for exoG, and doubly Hall for exoC. However in the critical case (III.10) the Galilean form (III.15) also decouples, and projecting into $2+1$ dimension, we get the same reduced form $\sigma_{exoC}$, (up to $B^{*}\leftrightarrow eB_{crit}$ and $s\leftrightarrow t$) — which is indeed identical to the DJT form $\sigma_{DJT}$ in (II.5) (up to $eB\leftrightarrow B^{*},\;t\leftrightarrow s,\;{\bf Q}\leftrightarrow{\bm{x}}\,.$ In conclusion :
Proposition IV.7.
Carrollian and critical Galilean motions project onto the same Hall motion, namely that for DJT in (II.2) (after identifying the guiding center ${\bf Q}$ with the DJT position).
@@ V Symmetry and immobility
V.1 Boost symmetry of DJT
This section is devoted to the curious double symmetry of the DJT model. We start with Galilei boosts,
${\bm{x}}\to{\bm{x}}+{\mathbf{b}}\,t\,,\quad t\to t\,,$ (V.1)
and observe that the lhs of the Hall law (II.2) is shifted, $(x_{i})^{\prime}\to(x_{i})^{\prime}+b_{i},$ whereas the rhs of (II.2) is left invariant. Thus :
@@ V.1 Boost symmetry of DJT
Galilei boost symmetry is broken for the DJT system as confirmed also in either the Lagrangian or in the symplectic framework :
${\mathscr{L}}_{DJT}$$\to$ ${\mathscr{L}}_{DJT}+\Big{(}-\hbox{$\textstyle\frac{1}{2}$}eB\epsilon_{ij}x_{i}b_{j}t+\hbox{$\textstyle\frac{1}{2}$}eE_{i}b_{i}t^{2}\Big{)}^{\prime}+eB\epsilon_{ij}x_{i}b_{j}\,,$ (V.2)
$\sigma_{DJT}$$\to$ $\sigma_{DJT}-eB\epsilon_{ij}b_{i}dx^{j}\wedge dt\,,$ (V.3)
@@
The unexpected result is that one of the Carroll boosts remains a symmetry : a C-boost (I.1), ${\bm{x}}\to{\bm{x}},\,s\to s-{\mathbf{b}}\cdot{\bm{x}}$, takes indeed the Souriau form (II.5) [up to trading Galilean time $t$ for Carroll time, $s$] into
$\sigma_{DJT}\to\sigma_{DJT}-e\epsilon_{ij}E_{i}b_{j}\,dx_{1}\wedge dx_{2}\,.$ (V.4)
The extra term vanishes when the boost is along the electric field. Viewed alternatively, the DJT Lagrangian changes as,
${\mathscr{L}}_{DJT}ds\to{\mathscr{L}}_{DJT}ds+d\Big{(}\hbox{$\textstyle\frac{1}{2}$}eE_{i}x_{i}^{2}b_{i}\Big{)}-e\Big{(}E_{1}b_{2}x_{1}dx_{2}+E_{2}b_{1}x_{2}dx_{1}\Big{)}\,,$ (V.5)
where the last term is a total derivative if $\epsilon_{ij}E_{i}b_{j}=0$. Thus, in contrast with the Galilean case :
@@
Conserved quantities are readily derived by the symplectic Noether theorem SSD recalled in Appendix A. Contracting the Souriau form (II.5) by an (infinitesimal) Carroll boost $\bm{\beta}$,
$X_{Carr}:\quad\delta{\bm{x}}=0{\quad\text{and}\quad}\delta s=-x_{i}\beta_{i}\,,$ (V.6)
@@
$\sigma(X\,,\,.\,)=eE\left\{\big{(}x_{1}dx_{2}\big{)}\beta_{1}+d\big{(}\frac{x_{2}^{2}}{2}\big{)}\beta_{2}\right\}\,.$ (V.7)
The 1st term here, obtained by boosting perpendicularly to ${\mathbf{E}}$ i.e. in the direction of the (Hall) motion, is not closed: $d\big{(}x_{1}dx_{2}\big{)}\beta_{1}=\big{(}dx_{1}\wedge dx_{2}\big{)}\beta_{1}\neq 0\,$. Thus :
@@
Boosting perpendicularly to ${\mathbf{E}}$ is not a symmetry, but boosting along ${\mathbf{E}}$ is a symmetry, with associated conserved quantity
${\mathcal{K}}_{||}=\frac{eE}{2}x_{2}^{2}{\quad\Rightarrow\quad}x_{2}=x_{2}^{0}\quad\text{fixed}\,,$ (V.8)
@@
The conservation of ${\mathcal{K}}_{||}$ can be checked also directly by using the equations of motion.
Rotational symmetry is broken when ${\mathbf{E}}\neq 0$, however the system is invariant w.r.t. space and Carroll-time translations,
$X_{C}=\gamma_{i}\frac{\partial}{\partial x_{i}}+\varepsilon\frac{\partial}{\partial s}\,,$ (V.9)
@@
$\begin{array}[]{lcll}p_{i}=&eB\epsilon_{ij}x_{j}+eE_{i}s&&\text{Carroll momentum}\\[4.0pt] {\mathscr{H}}=&eE_{i}x_{i}=eV({\bm{x}})&&\text{Carroll energy}\end{array}\quad.$ (V.10)
@@
These results could also by obtained in Hamiltonian terms. For example, $f_{2}$ Poisson-commutes with ${\mathscr{H}}$, $\big{\{}f_{2},{\mathscr{H}}\big{\}}=0$. Conversely, $f_{2}$ generates the Hamiltonian vectorfield $X_{1}={E}/{B},\;X_{2}=0\,.$ More generally, if $f=f(x_{1},x_{2})$ then $\{f,{\mathscr{H}}\}=\frac{1}{B}{\partial}_{1}f,$ which is conserved for an arbitrary function of $x_{2}$, $f=f(x_{2})$.
It is instructive to compare the Carroll results with those in the Galilean case: the Galilei boost symmetry is fully broken by the magnetic field, as we have seen. Accordingly, inserting the Galilean generators into the Souriau form yields,
$\sigma_{DJT}\big{(}X\,,\,\cdot\,\big{)}=d\Big{\{}\gamma_{i}e\big{(}B\varepsilon_{ij}x_{j}+E_{i}t\big{)}-\epsilon\big{(}eE_{i}x_{i}\big{)}+\beta_{i}\big{(}\frac{1}{2}eE_{i}t^{2}\big{)}\Big{\}}+\beta_{i}eBt\,,\epsilon_{ij}dx_{j}\,:\quad$ (V.11)

### v3

@@ IV Exotic Carroll particle
$B=0\,.$ (IV.13)
Then, assuming that $\,\theta\,\kappa_{mag}\neq 1,$ the position satisfies an anomalous Hall Law,
$\dot{x}_{i}=-\frac{e\theta}{1-\theta\kappa_{mag}}\epsilon_{ij}E_{j}\,,$ (IV.14)
even for $\kappa_{mag}=0$ as long as $\theta\neq 0$.
Such motions are reminiscent of “the” anomalous Hall effect observed in ferromagnetic materials in the absence of an external magnetic field Kar+ ; Johnson:1950zz ; Ezawa ; Stone ; AnomHall . Comparison of (IV.14) with (IV.7) underlines that $\kappa_{mag}$ behaves indeed as a proper magnetic field carried by the particle, and could be viewed as the planar analog of a magnetic monopole. It is also reminiscent of the rather mysterious magnetic field of relativistic anyons FeherAny .
IV.1 Chiral decomposition of exotic Carroll dynamics
Further insight is provided by chiral decomposition proposed in PlyushChiral1 ; PlyushChiral2 . Starting with the exotic Souriau form $\sigma_{exoC}$ (IV.2) [or symplectic form $\Omega_{exo}$ (III.3b) and Hamiltonian, ${\mathscr{H}}_{Carr}$ (IV.3)], we introduce chiral coordinates on the phase space,
$p_{i}=B^{*}\epsilon_{ij}X^{-}_{j}\,,\qquad x_{i}=X^{+}_{i}+X^{-}_{i}\,.$ (IV.15)
$X^{+}_{i}$ will later be renamed $Q_{i}$ and identified with the guiding centre Ezawa ; Stone . The symplectic form and Hamiltonian are decomposed as,
$\Omega_{exo}=\,$ $\Omega^{+}+\Omega^{-}\;\;\;=\Big{\{}\frac{B^{*}}{2}\epsilon_{ij}dX^{+}_{i}\wedge dX^{+}_{j}\Big{\}}-\Big{\{}\frac{B^{*}}{2}(1-{\theta}B^{*})\epsilon_{ij}dX^{-}_{i}\wedge dX^{-}_{j}\Big{\}},$ (IV.16a)
${\mathscr{H}}_{Carr}=\,$ ${\mathscr{H}}^{+}_{Carr}+{\mathscr{H}}^{-}_{Carr}=-eE_{i}X^{+}_{i}-eE_{i}X^{-}_{i}\,.$ (IV.16b)
We note the relative minus in (IV.16a) but not in (IV.16b).
$\left\{X^{+}_{i},X^{+}_{j}\right\}=-\frac{1}{B^{*}}\epsilon_{ij},\text{ \ \ }\left\{X^{-}_{i},X^{+}_{j}\right\}=0,\text{ \ \ }\left\{X^{-}_{i},X^{-}_{j}\right\}=\frac{1}{\left(1-{\theta}B^{*}\right)}\frac{1}{B^{*}}\epsilon_{ij}\,,$ (IV.17)
yield the Hamilton equations,
$\frac{dX^{+}_{i}}{ds}=\quad\;\epsilon_{ij}\frac{eE_{j}}{B^{*}}\,,\quad$ (IV.18a)
$\frac{dX^{-}_{i}}{ds}=\;-\frac{1}{\left(1-\theta{B^{*}}\right)}\epsilon_{ij}\frac{eE_{j}}{B^{*}}\,.$ (IV.18b)
Thus :
@@ IV.1 Chiral decomposition of exotic Carroll dynamics
Both chiral components $X^{+}_{i}$ and $X^{-}_{i}$ move along the same straight line, however their velocities which may differ,
${m^{*}}\,\frac{dX^{-}_{i}}{ds}=\;-\frac{dX^{+}_{i}}{ds}\,.$ (IV.19)
For $\theta=0$ the motions of the two components cancel and we recover the seminal immobility property of Carroll particles,
${\bm{x}}(s)={\bm{X}}^{+}(s)+{\bm{X}}^{-}(s)={\bm{x}}_{0}=\mathop{\rm const.}\nolimits$ (IV.20)
[…]
Projecting the 4-dimensional chiral phase space into 2-dimensional one the with the coordinate $X^{+}_{i}$ identified as the renamed Carroll guiding center,
$Q_{i}\equiv X^{+}_{i}=x_{i}+\frac{1}{B^{*}}\,\epsilon_{ij}p_{j}$ (IV.21)
cf. (III.12) we get, in all cases (and not only for $B_{crit}$),
$\Omega^{red}=\frac{1}{2}B^{*}\epsilon_{ij}dQ_{i}\wedge dQ_{j}\,,\qquad\big{\{}Q_{1},Q_{2}\big{\}}=-\frac{1}{B^{*}}\,,\qquad{\mathscr{H}}^{red}=-eE_{i}Q_{i}\,$ (IV.22)
where we recognize, putting ${\bf Q}\leadsto{\bm{\xi}},\,B^{*}\leadsto eB,\,s\leadsto t\,$ once again the DJT system of sec. II.
Exotic Galilean case. We briefly recall for comparison the exotic Galilean model described in chiral terms ZHChiral . The symplectic form and thus commutation relations are still (IV.16a) and (IV.17) (up to $B^{*}\leadsto eB$), but the Hamiltonian has a kinetic term which comes from ${\bm{p}}^{2}/2m$,
${\mathscr{H}}_{exoG}=\Big{\{}-e{\mathbf{E}}\cdot{\bm{X}}^{+}\Big{\}}\;+\;\Big{\{}\dfrac{e^{2}B^{2}}{2m}({\bm{X}}^{-})^{2}-e{\mathbf{E}}\cdot{\bm{X}}^{-}\Big{\}}\,.$ (IV.23)
The system splits into two uncoupled parts, however the $X_{i}^{-}$ equation has a kinetic term, cf. (IV.18).
$\begin{array}[]{rlr}\displaystyle\frac{dX^{+}_{i}}{dt}&=&\;\;\epsilon_{ij}\displaystyle\frac{E_{j}}{B}\,,\\[14.0pt] \left(1-\theta{eB}\right)\,\displaystyle\frac{dX^{-}_{i}}{dt}&=&\;eB\epsilon_{ij}X^{-}_{j}-\epsilon_{ij}\displaystyle\frac{E_{j}}{B}\,.\end{array}$ (IV.24)
For $m^{*}\neq 0$ the Galilean guiding centre follows the Hall law and $X_{i}^{-}$ rotates around it. In the critical case $m^{*}=0$, i.e. for $B=B_{crit}$ in (III.10) $X_{i}^{+}$ keeps on following the Hall law and the consistency of the 2nd eqn requires $X_{i}^{-}$ to be frozen into $X_{i}^{-}=E_{i}/eB^{2}$, consistently with the results in sec.III. In conclusion :
@@
Both Carrollian and critical Galilean motions project onto Hall motion, namely that for DJT in (II.2).
@@ V Symmetry and immobility
V.1 Boost symmetries of DJT
We first study the curious double symmetry of the DJT model.
$\bullet$ We start with Galilei boosts,
${\bm{\xi}}\to{\bm{\xi}}+{\mathbf{b}}\,t\,,\quad t\to t\,,$ (V.1)
and observe that the lhs of the Hall law (II.2) is shifted, $({\xi}_{i})^{\prime}\to({\xi}_{i})^{\prime}+b_{i},$ whereas the rhs of (II.2) is left invariant. Thus :
@@ V.1 Boost symmetries of DJT
Galilei boost symmetry is broken for the DJT system by the electromagnetic field.
This statement is readily confirmed also in either the Lagrangian or in the symplectic framework :
${\mathscr{L}}_{DJT}$ $\to{\mathscr{L}}_{DJT}+\Big{(}-\hbox{$\textstyle\frac{1}{2}$}eB\epsilon_{ij}{\xi}_{i}b_{j}t+\hbox{$\textstyle\frac{1}{2}$}eE_{i}b_{i}t^{2}\Big{)}^{\prime}+eB\epsilon_{ij}{\xi}_{i}b_{j}\,,$ (V.2a)
$\sigma_{DJT}$ $\to\sigma_{DJT}-eB\epsilon_{ij}b_{i}d{\xi}_{j}\wedge dt\,,$ (V.2b)
@@
$\bullet$ The unexpected result is that Carroll boosts, (I.1), behave somewhat better as their Galilei counterparts : a C-boost ${\bm{\xi}}\to{{\bm{\xi}}},\,s\to s-{\mathbf{b}}\cdot{{\bm{\xi}}}$ in (I.1) takes the Souriau form (II.5) [up to trading Galilean time $t$ for Carroll time, $s$] into
$\sigma_{DJT}\to\sigma_{DJT}-e\epsilon_{ij}E_{i}b_{j}\,d{\xi}_{1}\wedge d{\xi}_{2}\,,$ (V.3)
and the extra term vanishes when the boost is directed along the electric field. Alternatively, the DJT Lagrangian changes as,
${\mathscr{L}}_{DJT}ds\to{\mathscr{L}}_{DJT}ds+d\Big{(}\hbox{$\textstyle\frac{1}{2}$}eE_{i}{\xi}_{i}^{2}b_{i}\Big{)}-e\Big{(}E_{1}b_{2}{\xi}_{1}d{\xi}_{2}+E_{2}b_{1}{\xi}_{2}d{\xi}_{1}\Big{)}\,,$ (V.4)
when the boost is along the electric field, $\epsilon_{ij}E_{i}b_{j}=0$. Thus, in contrast with the Galilean case :
@@
Conserved quantities are conveniently derived by the symplectic Noether theorem SSD recalled in Appendix A. Contracting the Souriau form (II.5) by an infinitesimal Carroll boost $\bm{\beta}$,
$X_{b}\equiv X_{b}^{Carr}:\quad\delta{\bm{\xi}}=0{\quad\text{and}\quad}\delta s=-{\xi}_{i}\beta_{i}\,,$ (V.5)
@@
$\sigma_{DJT}(X_{b}\,,\,.\,)=eE\left(\big{(}{\xi}_{\perp}d{\xi}_{2}\big{)}\beta_{\perp}+d\big{(}\frac{{\xi}_{||}^{2}}{2}\big{)}\beta_{||}\right)\,.$ (V.6)
The first term here, obtained by boosting perpendicularly to ${\mathbf{E}}$ i.e. in the direction of the (Hall) motion, is not closed, $d\big{(}{\xi}_{\perp}d{\xi}_{||}\big{)}=\big{(}d{\xi}_{\perp}\wedge d{\xi}_{||}\big{)}\neq 0\,,$ but the second is closed. Thus :
@@
Boosting along ${\mathbf{E}}$ generates the conserved quantity
${\mathcal{K}}_{||}=\frac{eE}{2}{\xi}_{||}^{2}{\quad\Rightarrow\quad}{\xi}_{||}={\xi}_{||}^{0}\quad\text{fixed}\,,$ (V.7)
@@
The conservation of ${\mathcal{K}}_{||}$ can be checked also directly by using the equations of motion and could also by obtained in Hamiltonian terms.
For comparison, inserting the Galilean generators $X\equiv X_{b}^{Gal}$ into the Souriau form yields,
$\sigma_{DJT}\big{(}X\,,\,\cdot\,\big{)}=d\Big{\{}\gamma_{i}e\big{(}B\varepsilon_{ij}{\xi}_{j}+E_{i}t\big{)}-\epsilon\big{(}eE_{i}{\xi}_{i}\big{)}+\beta_{i}\big{(}\frac{1}{2}eE_{i}t^{2}\big{)}\Big{\}}+\beta_{i}eBt\,\epsilon_{ij}d{\xi}_{j}\,.\quad$ (V.8)
The system is invariant w.r.t. space and Carroll-time translations,
$X_{C}=\gamma_{i}\frac{\partial}{\partial{\xi}_{i}}+\varepsilon\frac{\partial}{\partial s}\,,$ (V.9)
@@
$\begin{array}[]{llclc}p_{i}&=&eB\epsilon_{ij}{\xi}_{j}+eE_{i}s&&\text{(Carroll) momentum}\\[4.0pt] {\mathscr{H}}_{Carr}&=&eE_{i}{\xi}_{i}&&\text{(Carroll) energy}\end{array}\quad.$ (V.10)
@@
Rotational symmetry is manifestly broken when ${\mathbf{E}}\neq 0$, as confirmed also by inserting the infinitesimal rotation generator $X_{r}:\delta\xi_{i}=\epsilon_{ij}\xi_{j},\,\delta s=0$ into the Lagrange form,
$\sigma_{DJT}\big{(}X_{r},\,\cdot\,\big{)}=d\big{(}{\hbox{$\textstyle\frac{1}{2}$}}eB|{\bm{\xi}}|^{2}\big{)}+\epsilon_{ij}eE_{i}\xi_{j}\,ds\,,$ (V.11)
where the last term is not a total differential.